OpenStudy (anonymous):

If I took 7 hours to mow 4 lawns, then at that rate, how many lawns could be mowed in 35 hours.?At what rate were lawns being mowed.?

OpenStudy (anonymous):

help

OpenStudy (anonymous):

d=r/t d is unknown therefore r is 4 lawns mown t is 35 hours 35-7=28 so d=28/4 d=7

OpenStudy (anonymous):

7 lawns were mown in 35 hours